a car atarts moving on a straight road with constant acceleration of 2m/sec2. at the same instant a truck moving at constant velocity of 20m/sec passes by the car find the time taken by the carto just overtake the truck.

Grade:9

1 Answers

swarup kumar majee
6 years ago
For the car,
u=0m/s,a=2m/s2
so,in t=1s it will cover a distance of 1m
and,the truck will cover 20m
so,the distance between the car and the truck will be 19m
if,in x distance the car will overtake the truck then,
x=2t+t2 or t2+2t-x=0           ---------------------1
x=20t+19                          -----------------------2
putting value of x in equation 1
t2-18t-19=0
t=18\pmroot400/2
t=18s+1s=19s(because before also we took that t=1)